Fermi National Laboratory
 

Accelerator Update for Jun 14, 2006 - June 15, 2006

 

Wednesday June 14

The day shift began with Operations monitoring store 4760 and a Recycler stash, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, with the Antiproton Source (Pbar) stacking, with the Booster, Main Injector (MI), and Recycler conducting studies, with NuMI off due to a problem with a Lambertson power supply (I:LAM61), and with MiniBooNE taking beam.

Operations prepared MI-13 for access at 8:39 AM.

Experts investigated LAM61 and reported at 9:51 AM that the regulator would have to be replaced. This repair will be done this afternoon. They made a temporary repair to the supply so that NuMI could resume running. Operations resumed sending them beam.

Operations began preparing the Recycler at 10:29 AM to stash antiprotons.

A Tevatron Electron Lens (TEL) expert completed adjustments to the TEL pulse width at 11:25 AM.

The stash ended and Pbar resumed stacking at 12:09 PM.

At 1:15 PM, the NuMI beam permit tripped off due to a Horn problem. The Horn reset okay.

At 1:46 PM, store 4760 aborted due to a TeV sector A0 coasting beam valve closure, but the valves never closed. Operators investigated and found a technician replacing a bad Ethernet module directly below the gate valve control unit, which is connected into the abort system. The controls experts investigated further and found loose connectors.

Operations halted stacking at 1:59 PM while experts replaced a Pbar Debuncher power supply (D:VT716). Stacking resumed at 2:13 PM.

Operations prepared D0 for access at 2:36 PM.

A MI expert reported at 5 PM that the MI-31 and MI-40 BLMs were not working.

D0 came out of access at 5:30 PM.

Recycler began preparing to stash at 5:35 PM.

At 5:51 PM, Operations reported that a Debuncher quadrupole power supply (D:Q916) was operating erratically. An expert was called in.

Operations put Pbar back to a stacking mode at 6:19 PM.

Operations put the TeV into a shot setup at 6:27 PM.

At 8:20 PM, while starting to ramp, the TeV suffered a sector B2 quench. Operators called in TeV and TEL experts to investigate.

Pbar resumed stacking around 9:20 PM.

A power supply expert completed work on D:Q916. Recycler resumed stashing at 8:54 PM. The stash ended at 9:21 PM.

At 10:37 PM, Operations reported a problem with the TeV B2 dump. The DC breaker wasn’t closing. Experts investigated.

Thursday June 15

The midnight shift began with Operations monitoring the stash, with Pbar stacking, with the TeV recovering from a B2 dump fault, with MI and Recycler conducting studies, and with MiniBooNE and NuMI taking beam.

Operations reported at 12:12 AM that experts replaced a faulty solenoid for the B2 DC dump breaker.

Operations turned on the TeV and TeV studiers began aperture scans at 1:22 AM. The scans ended 3:27 AM.

Operations put the TeV into a wet squeeze at 4:04 AM.

The Recycler began preparing to stash at 4:35 AM. The stash ended at 5:21 AM.

Operations put the TeV into a dry squeeze at 5:42 AM.

Operations put the TeV into a shot setup at 7:09 AM.

At 7:49 AM, a Linac RF station (KRF6) tripped off and wouldn’t reset. Operators investigated.

The Plans for Thursday

The plans for today are to establish a new store and to push proton and antiproton intensities.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/16/2006    email Fermilab